# Scalable Feature Learning on Huge Knowledge Graphs for Downstream Machine Learning

### v1

@@ C.4 Evaluation on knowledge graph completion
Nevertheless, Figure 8 shows that a Friedman test followed by Conover’s post-hoc analysis [Conover and Iman, 1979, Conover, 1999] reveal no statistically significant differences (at significance level $\alpha=0.05$) among the three methods that scale to all the knowledge graphs (SEPAL, PBG, and DGL-KE), as indicated by the critical difference diagrams where all methods are connected by a black line.
@@
Downstream and link prediction performance weakly correlate
@@ D.1 Analysis of SEPAL’s dynamic and analogies to eigenvalue problems
followed by blockwise normalization of $\boldsymbol{y}^{(t)}$ (with a $\ell_{2,\infty}$ mixed norm), shares structural similarities with the Arnoldi iteration [Arnoldi, 1951]:
@@
The Arnoldi iteration is used to compute numerical approximations of the eigenvectors of general matrices, for instance, in ARPACK [Lehoucq et al., 1998]. This analogy suggests that the direction of outer embeddings stabilizes over time and aligns with a form of dominant generalized eigenvector of the propagation operator $\boldsymbol{M}$, conditioned on the core. Therefore, the embeddings produced by SEPAL’s propagation encapsulate global structural information on the knowledge graph.
D.2 Formal proof of Proposition 4.1
@@
Therefore, a parallel can be drawn between: 1) the outer subgraphs and mini batches, 2) the number of propagation steps $T$ and number of epochs, 3) the hyperparameter $\alpha$ and the learning rate.
After each gradient step, we normalize the entity embeddings to enforce the unit norm constraint. This procedure corresponds to projected gradient descent on the sphere [Bertsekas, 1997], where each update is followed by a projection (via $\ell^{2}$ normalization) back onto the feasible set. The energy function $\mathcal{E}$ (Equation 5) is composed of inner products and element-wise multiplications of smooth functions, thus it is smooth, and its gradient is Lipschitz continuous on the unit sphere. Under these conditions, the algorithm is guaranteed to converge to a stationary point of the constrained optimization problem [Bertsekas, 1997, Proposition 2.3.2]. The limit points of the optimization thus satisfy the first-order optimality conditions on the sphere.
@@
  $\mathbb{S}\leftarrow\emptyset$triangleright list of subgraphs
  $U\leftarrow V$triangleright set of unassigned nodes
@@
  while $|U|>(1-h)|V|$ do
     $k\leftarrow 0$  ;  $\mathcal{S}_{0}\leftarrow\{\operatorname*{arg\,max}_{v\in U}deg(v)\}$triangleright start with unassigned node $v$ with highest degree
@@
     Append $\mathcal{S}_{k-1}$ to $\mathbb{S}$, and update $U$triangleright $\mathcal{S}_{k-1}$ is the last subgraph smaller than $0.8\,m$
@@ E.1 Prior work on graph partitioning
Scaling up computation on graph, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. METIS [Karypis and Kumar, 1997], a greedy node-merging algorithm, is a popular solution. A variety of algorithms have also been developed to detect “communities”, groups of nodes more connected together, often with applications on social networks: Spectral Clustering (SC) [Shi and Malik, 2000], the Leading Eigenvector (LE) method [Newman, 2006], the Label Propagation Algorithm (LPA) [Raghavan et al., 2007], the Louvain method [Blondel et al., 2008], the Infomap method [Rosvall and Bergstrom, 2008], and the Leiden method [Traag et al., 2019] which guarantees connected communities. LDG [Stanton and Kliot, 2012] and FENNEL [Tsourakakis et al., 2014] are streaming algorithms for very large graphs. Some algorithms have also been specifically tailored for power-law graphs: DBH [Xie et al., 2014] leverages the skewed degree distributions to reduce the communication costs, HDRF [Petroni et al., 2015] is a streaming partitioning method that replicates high-degree nodes first, and Ginger [Chen et al., 2019b] is a hybrid-cut algorithm that uses heuristics for more efficient partitioning on skewed graphs.
@@ E.2 Detailed algorithm and pseudocode
High-level principle
BLOCS uses three base mechanisms to grow the subgraphs: diffuse (add all neighboring entities to the current subgraph), merge (merge two overlapping subgraphs), and dilate (add all unassigned neighboring entities to the current subgraph). There are two different regimes during the generation of subgraphs. First, few entities are assigned, and the computationally effective diffusion quickly covers a large part of the graph, especially entities that are close to high-degree nodes. However, once these close entities have been assigned, the effectiveness of diffusion drops because it struggles to reach entities farther away. For this reason, BLOCS switches from diffusion to dilation once the proportion of assigned entities reaches a certain threshold $h$ (a hyperparameter chosen .$\approx.6$, depending on the dataset). By adding only unassigned neighbors to subgraphs, dilation drives subgraph growth towards unassigned distant entities. However, the presence of long chains can drastically slow down this regime because they make it add entities one by one. Some knowledge graphs have long chains, for instance YAGO4.5 (see Diameter in Table 2). To tackle them, BLOCS switches back to diffusion for a few steps, with seeds taken inside the long chains.
BLOCS works faster on graphs that have small diameters, where most entities can be reached during the diffusion regime and fewer dilation steps are required (Section F.1).
Figure 10: Scalability of partitioning methods. Memory usage and time for 8 knowledge graphs of various sizes. The “CPU RAM capacity” dashed line represents the CPU RAM of the machine we used to run SEPAL (we used a different machine with more RAM to run this benchmark, see Section B.3). Partitioning methods going beyond this limit thus cannot be combined with SEPAL on our hardware. BLOCS is the only method to scale up to WikiKG90Mv2, a knowledge graph with 601M triples. Leiden and METIS both caused memory errors on WikiKG90Mv2, while HDRF was too long for graphs larger than YAGO4 (our time limit was set to 333 minutes).
@@
Detailed description
Algorithm 1 describes BLOCS pseudocode. Below, we provide more details on subparts of the algorithm.
@@
The function SystematicMerge is very similar to MergeSmallSubgraphs, but it does not check that the resulting subgraphs are smaller than $m$. Its objective is to eliminate all the subgraphs whose size is smaller than a given threshold (set to $0.4\,m$ in Algorithm 1). The subgraphs produced that are larger than $m$ are then handled by the function SplitLargeSubgraphs.
@@ E.3 Benchmarking BLOCS against partitioning methods
is based on a multilevel recursive bisection approach, which coarsens the graph, partitions it, and then refines the partitions. It produces partitions with the same sizes; however, due to the graph structure, they often explode into multiple connected components, which is detrimental to the embedding propagation (see Section E.3).
@@ F.2 SEPAL’s hyperparameters
List of SEPAL’s hyperparameters
@@
Experimental study of hyperparameter effect
@@
Strategy $\boldsymbol{\eta_{n}}$ $\boldsymbol{\eta_{e}}$Rel #Entities #Triples Time
@@
    It yields very dense core subgraphs. Indeed, while they contain $\eta_{n}$of the entities of the full graph, they gather around $4\eta_{n}$of all the triples (Table 8). This allows the training on the core to process a substantial portion of the knowledge-graph triples, resulting in richer representations.
@@
Regarding performance, Figure 20 shows that the hybrid strategy slightly underperforms the degree-based approach on four real-world downstream tasks. Indeed, the hybrid approach enhances relational coverage but, as a counterpart, yields a sparser core subgraph (see #Triples in Table 8), which is detrimental to downstream performance. However, one can adjust the values of $\eta_{n}$ and $\eta_{e}$ to control the trade-off between downstream performance and relation coverage, depending on the use case. For instance, for the link prediction task, a better relational coverage (greater $\eta_{e}$) improves the performance (Section C.4).
@@ F.4 Speedup over base embedding model
Figure 22 demonstrates that SEPAL can accelerate its base embedding model by more than a factor of 20, while also its boosting performance on downstream tasks.
@@ Appendix G Discussion
G.1 Comparison between SEPAL and NodePiece
Figure 23: Comparing SEPAL with NodePiece on Mini YAGO3.
@@ G.2 Communication costs of SEPAL
In modern computing architectures, memory transfer costs are high, amounting to much computation, and the key to achieve high operation efficiency is to reduce data movement [Mutlu et al., 2022].
@@
Dataset $x_{\mathrm{outer}}$
@@
G.3 Broader impacts
SEPAL may reflect the biases present in the training data. For instance, Wikipedia, from which YAGO is derived, under-represents women [Reagle and Rhue, 2011]. We did not evaluate how much our method captures such biases. We note that the abstract nature of embeddings may make the biases less apparent to the user; however, this problem is related to embeddings and not specific to our method. It may be addressed by debiasing techniques [Bolukbasi et al., 2016, Fisher et al., 2020] for which SEPAL could be adapted.

### v2

@@ C.6.1 Link prediction results
Nevertheless, Figure 8 shows that a Friedman test followed by Conover’s post-hoc analysis [conover1979multiple, conover1999practical] reveal no statistically significant differences (at significance level $\alpha=0.05$) among the three methods that scale to all the knowledge graphs (SEPAL, PBG, and DGL-KE), as indicated by the critical difference diagrams where all methods are connected by a black line.
@@
C.6.2 Downstream and link prediction performance weakly correlate
@@ D.1 Analysis of SEPAL’s dynamic and analogies to eigenvalue problems
followed by blockwise normalization of $\boldsymbol{y}^{(t)}$ (with a $\ell_{2,\infty}$ mixed norm), shares structural similarities with the Arnoldi iteration [arnoldi1951principle]:
@@
The Arnoldi iteration is used to compute numerical approximations of the eigenvectors of general matrices, for instance, in ARPACK [lehoucq1998arpack]. This analogy suggests that the direction of outer embeddings stabilizes over time and aligns with a form of dominant generalized eigenvector of the propagation operator $\boldsymbol{M}$, conditioned on the core. Therefore, the embeddings produced by SEPAL’s propagation encapsulate global structural information on the knowledge graph.
D.2 Formal proof of Proposition˜4.1
@@
Therefore, a parallel can be drawn between: 1) the outer subgraphs and mini-batches, 2) the number of propagation steps $T$ and the number of epochs, 3) the hyperparameter $\alpha$ and the learning rate.
After each gradient step, we normalize the entity embeddings to enforce the unit norm constraint. This procedure corresponds to projected gradient descent on the sphere [bertsekas1997nonlinear], where each update is followed by a projection (via $\ell^{2}$ normalization) back onto the feasible set. The energy function $\mathcal{E}$ (Equation 7) is composed of inner products and element-wise multiplications of smooth functions, thus it is smooth, and its gradient is Lipschitz continuous on the unit sphere. Under these conditions, the algorithm is guaranteed to converge to a stationary point of the constrained optimization problem [bertsekas1997nonlinear, Proposition 2.3.2]. The limit points of the optimization thus satisfy the first-order optimality conditions on the sphere.
@@
  $\mathbb{S}\leftarrow\emptyset$ $\triangleright$ list of subgraphs
  $U\leftarrow V$ $\triangleright$ set of unassigned nodes
@@
  while |U|>(1h)|V||U|>(1-h)|V| do
   $k\leftarrow 0$    ;    $\mathcal{S}_{0}\leftarrow\{\operatorname*{arg\,max}_{v\in U}deg(v)\}$ $\triangleright$ start with unassigned node $v$ with highest degree
@@
   Append $\mathcal{S}_{k-1}$ to $\mathbb{S}$, and update $U$ $\triangleright$ $\mathcal{S}_{k-1}$ is the last subgraph smaller than 0.8m0.8\,m
@@ E.1 Prior work on graph partitioning
Scaling up computation on graph, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. METIS [karypis1997metis], a greedy node-merging algorithm, is a popular solution. A variety of algorithms have also been developed to detect “communities”, groups of nodes more connected together, often with applications on social networks: Spectral Clustering (SC) [shi2000normalized], the Leading Eigenvector (LE) method [newman2006finding], the Label Propagation Algorithm (LPA) [raghavan2007near], the Louvain method [blondel2008fast], the Infomap method [rosvall2008maps], and the Leiden method [traag2019louvain] which guarantees connected communities. LDG [stanton2012streaming] and FENNEL [tsourakakis2014fennel] are streaming algorithms for very large graphs. Some algorithms have also been specifically tailored for power-law graphs: DBH [xie2014distributed] leverages the skewed degree distributions to reduce the communication costs, HDRF [petroni2015hdrf] is a streaming partitioning method that replicates high-degree nodes first, and Ginger [chen2019powerlyra] is a hybrid-cut algorithm that uses heuristics for more efficient partitioning on skewed graphs.
@@ E.2 Detailed algorithm and pseudocode
Figure 10: Scalability of partitioning methods. Memory usage and time for 8 knowledge graphs of various sizes. The “CPU RAM capacity” dashed line represents the CPU RAM of the machine we used to run SEPAL (we used a different machine with more RAM to run this benchmark, see Section˜B.3). Partitioning methods going beyond this limit thus cannot be combined with SEPAL on our hardware. BLOCS is the only method to scale up to WikiKG90Mv2, a knowledge graph with 601M triples. Leiden and METIS both caused memory errors on WikiKG90Mv2, while HDRF was too long for graphs larger than YAGO4 (our time limit was set to 333 minutes).
@@
Algorithm˜1 describes BLOCS pseudocode. Below, we provide more details on subparts of the algorithm.
@@
The function SystematicMerge is very similar to MergeSmallSubgraphs, but it does not check that the resulting subgraphs are smaller than $m$. Its objective is to eliminate all the subgraphs whose size is smaller than a given threshold (set to 0.4m0.4\,m in Algorithm˜1). The subgraphs produced that are larger than $m$ are then handled by the function SplitLargeSubgraphs.
@@ E.3 Benchmarking BLOCS against partitioning methods
is based on a multilevel recursive bisection approach, which coarsens the graph, partitions it, and then refines the partitions. It produces partitions with the same sizes; however, due to the graph structure, they often explode into multiple connected components, which is detrimental to the embedding propagation (see Section˜E.3).
@@ F.2 SEPAL’s hyperparameters
F.2.1 List of SEPAL’s hyperparameters
@@
F.2.2 Experimental study of hyperparameter effect
@@
Strategy $\boldsymbol{\eta_{n}}$ $\boldsymbol{\eta_{e}}$ #Rel #Entities #Triples Time
@@
    It yields very dense core subgraphs. Indeed, while they contain $\eta_{n}$% of the entities of the full graph, they gather around $4\eta_{n}$% of all the triples (Table 10). This allows the training on the core to process a substantial portion of the knowledge-graph triples, resulting in richer representations.
@@
Regarding performance, Figure 20 shows that the hybrid strategy slightly underperforms the degree-based approach on four real-world downstream tasks. Indeed, the hybrid approach enhances relational coverage but, as a counterpart, yields a sparser core subgraph (see #Triples in Table 10), which is detrimental to downstream performance. However, one can adjust the values of $\eta_{n}$ and $\eta_{e}$ to control the trade-off between downstream performance and relation coverage, depending on the use case. For instance, for the link prediction task, a better relational coverage (greater $\eta_{e}$) improves the performance (Section˜C.6).
@@
Table 11: Effect of core selection strategy: PageRank vs Degree. We compare the performance of SEPAL+DistMult on real-world downstream tasks when using either degree or PageRank as a centrality measure for core selection. We report the average normalized R2 score across the four downstream tasks, along with the total execution time (in parentheses).
Dataset Core defined by Degree Core defined by PageRank
YAGO3 0.783 (11m 20s) 0.742 (9m 23s)
YAGO4 0.817 (6h 20m) 0.861 (4h 53m)
YAGO4+T 0.815 (10h 9m) 0.881 (8h 16m)
YAGO4.5 0.949 (4h 11m) 0.925 (4h 3m)
YAGO4.5+T 0.923 (2h 47m) 0.912 (2h 40m)
Freebase 0.917 (5h 58m) 0.919 (5h 58m)
WikiKG90Mv2 0.898 (20h 31m) 0.902 (23h 59m)
Exploring other centrality measures
Here, we compare degree with PageRank as a centrality measure for the core selection (keeping all other hyperparameters unchanged). Like the degree, the PageRank can be computed very efficiently on huge graphs using sparse matrix multiplications.
Table 11 reports results on the real-world downstream tasks. Regarding performance, it seems that the difference between degree and PageRank depends on the graph’s structure. Specifically, we observe that for YAGO4, YAGO4+T, Freebase, and WikiKG90Mv2, PageRank improves the performance. This echoes the results of Figure 14, which show that in these four specific KGs, contrary to the others, the degree-based core selection yields cores that are connected through fewer edges to some outer subgraphs. We can therefore conjecture that for these KGs, PageRank improves information flow and mitigates issues like oversquashing during propagation, ultimately increasing performance. Regarding computational cost, we see that SEPAL with PageRank usually runs slightly faster than with degree. This is because PageRank yields slightly sparser cores, leading to faster core training.
@@ F.4 Speedup over base embedding model
Figure 22 demonstrates that SEPAL can accelerate its base embedding model by more than a factor of 20, while also boosting its performance on downstream tasks.
@@ Appendix G Discussion
G.1 Comparison to prior work
G.1.1 Comparison to DistMult-ERAvg
albooyeh2020out propose an aggregation that is similar to SEPAL’s aggregation during the propagation phase, however the two methods optimize the embeddings in two very distinct ways:
  • albooyeh2020out introduce propagation within the standard link prediction pipeline: during training, for each triple $(v,r,u)$, they occasionally (with probability $p$) replace one entity’s embedding with an aggregated version (e.g., $\theta_{u}\cdot\theta_{r}$) before computing the plausibility score. However, training still relies on negative sampling and a classic link prediction loss, and thus optimizes for local triple-level contrasts like traditional KGE methods.
  • SEPAL, in contrast, explicitly separates the optimization objective: embeddings for a small core are trained with a classic KGE objective, and then relation-aware propagation is used across the rest of the graph, without negative sampling. This distinction is crucial: SEPAL removes the need for negative sampling on typically 95 to 99% of the graph, enabling both improved scalability and alignment properties beneficial for downstream tasks.
Moreover, DistMult-ERAvg is not designed for very large graphs. On the datasets considered in this paper, it fails with out-of-memory errors on all the graphs except Mini YAGO3 (129k entities, 1.1M triples).
Table 12: Comparison between SEPAL and DistMult-ERAvg on Mini YAGO3. We report the normalized mean cross-validation score (R2) across the four real-world downstream tasks, along with the total training time. SEPAL outperforms DistMult-ERAvg while being significantly faster.
Method Housing prices Movie revenues US accidents US elections Time
DistMult-ERAvg 0.149 0.124 0.444 0.916 2h 41m
SEPAL 0.276 0.159 0.548 0.929 5m
Table 12 reports performance and runtime on Mini YAGO3, showing that SEPAL is 32 times faster than DistMult-ERAvg while achieving consistently better scores. This is expected since DistMult-ERAvg follows the classic optimization loop with negative sampling, gradient computations, and parameter updates, and therefore inherits the limitations of traditional KGE methods. The strength of DistMult-ERAvg lies in out-of-sample embedding computation, which SEPAL also supports (see Section˜G.3), but with greater scalability.
G.1.2 Comparison to NodePiece
@@
Figure 23: Comparing SEPAL with NodePiece on Mini YAGO3.
@@ G.2 Communication costs of SEPAL
In modern computing architectures, memory transfer costs are high, amounting to much computation, and the key to achieve high operation efficiency is to reduce data movement [mutlu2022modern].
@@
Dataset $\boldsymbol{x_{\mathrm{outer}}}$
@@
G.3 Outlook on continual learning
Table 14: Statistics of the two versions of YAGO3 used to illustrate SEPAL’s suitability for continual learning.
Version #Entities #Relations #Triples
YAGO3-2014 2,570,716 37 5,585,004
YAGO3-2022 4,546,966 37 14,691,781
The modular nature of SEPAL makes it well-suited for continual learning scenarios, where new entities are added to the knowledge graph over time. Indeed, new embeddings can be computed without retraining from scratch, via a few additional propagation steps, as long as the relations remain unchanged. To demonstrate this, we use two versions of the YAGO3 knowledge graph: the original 2014 release, and the 2022 revived version. Table 14 gives the statistics of these two datasets.
Table 15: Continual learning experiment on YAGO3. Average normalized R2 scores across the four real-world downstream regression tasks (movie revenues, US accidents, US elections, housing prices) and total runtime for different methods. SEPAL-CL denotes SEPAL in the continual learning setting, where new entities are embedded via propagation.
YAGO3 version Method Average performance Runtime
2014 SEPAL 0.836 0h 11m 20s
2022 DistMult 0.884 11h 05m 36s
2022 SEPAL 0.988 1h 05m 07s
2022 SEPAL-CL 0.962 0h 01m 08s
We adapt SEPAL to this continual learning setting by: (1) initializing the embeddings of YAGO3-2022 using embeddings precomputed on YAGO3-2014, (2) propagating embeddings for 5 additional steps to update existing nodes and embed new entities. We denote this method SEPAL-CL in the results.
Table 15 shows that:
  • For downstream applications, YAGO3-2022 brings value compared to YAGO3-2014. Indeed, for all the methods considered, the embeddings learned on the 2022 dataset score higher than the strongest method on YAGO3-2014, SEPAL.
  • SEPAL trained from scratch is 10 $\times$ faster than DistMult on YAGO3-2022.
  • SEPAL-CL is 57 $\times$ faster than SEPAL trained from scratch on YAGO3-2022, and 587 $\times$ faster than DistMult.
  • SEPAL-CL clearly outperforms DistMult, and almost matches the performance of SEPAL trained from scratch, even in this very challenging scenario (8 years between the two versions of the graph), where the size of the graph has doubled in terms of entities, and tripled in terms of triples. In a real-life application, we could imagine embeddings recomputed monthly at very low cost.
G.4 Broader impacts
SEPAL may reflect the biases present in the training data. For instance, Wikipedia, from which YAGO is derived, under-represents women [reagle2011gender]. We did not evaluate how much our method captures such biases. We note that the abstract nature of embeddings may make the biases less apparent to the user; however, this problem is related to embeddings and not specific to our method. It may be addressed by debiasing techniques [bolukbasi2016man, fisher2020debiasing] for which SEPAL could be adapted.